# JASA-supplement

### v3

@@ 7.2 Sampling TMVN over regions of interest
Once parameters have been estimated, we can sample the censored data given the observed data as in Section 6.5, holding the first $n_{1}$ observed data fixed. Unlike for parameter estimation, we recommend using variable reordering for sampling TMVN distributions to improve the acceptance rate. However, the acceptance rate for sampling the TMVN distribution can be very low even with exponential tilting when the number of censored responses is high (e.g., $(n-n_{1})>1{,}000$), making it infeasible to obtain a reasonable number of samples for posterior inference. In this case, we propose to divide the censored locations according to subregions of the input space, each containing a manageable number (say, at most $1{,}000$) of censored locations; then we can draw TMVN samples over each region of interest separately. The responses at censored locations over one subregion are assumed to be conditionally independent from those over another subregion, given all observed responses. This idea is intuitive and simple to implement. Its downside is that it ignores censored locations outside of the selected region. As a heuristic remedy, we extend the boundary of the region of interest when drawing TMVN samples to incorporate surrounding information; this extension reduces the bias but also the acceptance rate. In many applications, obtaining the joint posterior distribution of the quantity of interest over a subregion is sufficient; for example, one could obtain the joint distribution of a water-related quantity over a catchment area and then compute the distribution of functions of interests (e.g., the average) as necessary.
To demonstrate this idea, we first considered an example in sufficiently low dimensions so that the ‘global’ simulation of all censored locations had a reasonable acceptance rate above $10^{-4}$. Specifically, we considered Scenario 1 in Table 1, where the censoring threshold is zero. Figure 7 compares $N=1{,}000$ samples of the north-west quarter (i.e., $[0,0.5]\times[0.5,1.0]$) generated by the global and regional sampling schemes using VMET.
@@
For the regional sampling scheme, we sampled the TMVN distribution defined over the region $[0,0.6]\times[0.4,1.0]$, which enveloped the region of our interest, conditional on all observed locations. For the global sampling scheme, we needed to sample from a $380$-dimensional TMVN distribution, whereas for the regional scheme, we only needed to sample from a $125$-dimensional TMVN distribution, resulting in a $50$ to $100$ times difference in acceptance rates. The two groups of samples were very similar except for minor differences in the tails, which can be attributed to the reduced number of censored locations used in the regional sampling scheme. Conditioning on more censored locations would decrease the conditional mean under our Matérn covariance kernel. We also compared the RMSE for predicting the censored data in the north-west quarter, where the global and regional sampling produced errors of $0.24$ and $0.25$, respectively. In contrast, the GP with the LOD-augmented data produced a prediction error of $0.62$ and the stochastic approximation of expectation maximization (SAEM), a state-of-the-art method for sampling TMVN distributions from Ordoñez et al. (2018), produced an error of $0.47$. For SAEM, we set the MC sample size to $500$ as in Ordoñez et al. (2018), resulting in a computation time of approximately two hours, which was significantly higher than for our proposed methods. The regional scheme had much higher accuracy than the competing methods and the minor accuracy loss compared with the global scheme, we believe, is a good trade-off for the gain in computational efficiency.
@@
Next, we considered an example in higher dimensions to demonstrate that the VMET method can be used in combination with the regional sampling scheme to improve posterior inference when there are censored GP responses. We again considered Scenario 1 from Table 1, but now with $n=6{,}400$. The region of interest and the region over which we drew TMVN samples were the same as in the low-dimensional example above. We had a total of $3{,}278$ censored locations, with $1{,}075$ inside our sampling region. $N=1{,}000$ samples were drawn and used in combination with all observed responses to predict the responses at $500$ randomly selected testing locations in our region of interest (i.e., $[0,0.5]\times[0.5,1.0]$). For comparison, we predicted with the GP using the LOD-augmented data; we also attempted a comparison to SAEM, but the approach was computationally infeasible for sampling the $1{,}075$-dimensional TMVN distribution due to extremely low acceptance rates. The LOD-augmented GP produced an RMSE of $0.45$, whereas the RMSE of our method was only $0.19$. While the global approach cannot be computed as a gold-standard comparison method in this high-dimensional example, note that the nugget variance is $0.03$, indicating a minimum prediction RMSE of $0.17$. This highlights the benefits of properly taking the censored locations into consideration for posterior inference. In this example, the acceptance rate was approximately $2\times 10^{-5}$ and the sampling took less than 44 minutes. Combined with the results from Section 7.1, the VMET algorithm outlined in Algorithm 1 enables scalable (linear-complexity) computation for the partially censored MVN model in terms of parameter estimation and posterior inference, significantly outperforming the LOD-augmented GP and the SAEM approaches.
@@ 7.3 Analysis of groundwater contamination
We considered groundwater tetrachloroethylene concentrations from the United States Geological Survey (USGS) to demonstrate that our proposed VMET method can be used in combination with the censored MVN model to enhance the modeling of partially censored data. Tetrachloroethylene can be used in dry cleaning, manufacturing other chemicals, and cleaning metals, but exposure to tetrachloroethylene is harmful and may cause skin and respiratory problems. The data were collected at irregular locations across the United States over the time period 2000 to 2022. In this section, the locations $\{\mathbf{s}_{i}\in\mathbb{R}^{3}\}$ are spatial-temporal coordinates consisting of longitudes, latitudes, and dates. Overall, there are $n=24{,}701$ spatio-temporal responses, among which there are $20{,}730$ censored responses due to detection limits. The detection thresholds are known and vary across spatio-temporal locations. Environmental contaminants are usually modeled using MVN distributions after a logarithm transformation (Helsel, 1990). During data preprocessing, we first applied a log-transform to both observations and detection thresholds. Next, we normalized the transformed observations and the detection thresholds with the mean and standard deviation of the transformed observations. The spatial and temporal coordinates were linearly scaled into the unit hypercube in $\mathbb{R}^{3}$, with the two spatial coordinates scaled by the same constant. The dataset, including observations and detection thresholds, is randomly split into 80% training and 20% testing.
@@
We used the state of Texas as our region of interest for posterior inference. Specifically, our goal was to make predictions over a dense regular $0.13^{\circ}\times 0.11^{\circ}$ longitude-latitude grid covering Texas on the last day of the dataset’s temporal range. Based on the predictions, we produced a ‘heatmap’ for the groundwater tetrachloroethylene concentrations in Texas. Conditional on the training observations, we drew $N=1{,}000$ samples of the TMVN distribution defined over the $546$ censored spatio-temporal points in the training dataset that are located in Texas. Inference over the grid as well as the testing spatio-temporal points in Texas was made by kriging using observed responses and the TMVN samples drawn previously. For comparison, we also considered predictions made by a GP using the augmented data where censored responses were replaced by their LODs. The SAEM method was computationally infeasible for sampling from the $546$-dimensional TMVN distribution, and it was hence excluded from our comparison. Figure 8 compares the heatmaps produced by the partially censored MVN model and the GP model.

### v4

@@ 7.2 Sampling TMVN over regions of interest
Once parameters have been estimated, we can sample the censored data given the observed data as in Section 6.5, holding the first $n_{1}$ observed data fixed. Unlike for parameter estimation, we recommend using variable reordering for sampling TMVN distributions to improve the acceptance rate. However, the acceptance rate for sampling the TMVN distribution can be very low even with exponential tilting when the number of censored responses is high (e.g., $(n-n_{1})>1{,}000$), making it infeasible to obtain a reasonable number of samples for posterior inference. In this case, we propose to divide the censored locations according to subregions of the input space, each containing a manageable number (say, at most 1,0001{,}000) of censored locations; then we can draw TMVN samples over each region of interest separately. The responses at censored locations over one subregion are assumed to be conditionally independent from those over another subregion, given all observed responses. This idea is intuitive and simple to implement. Its downside is that it ignores censored locations outside of the selected region. As a heuristic remedy, we extend the boundary of the region of interest when drawing TMVN samples to incorporate surrounding information; this extension reduces the bias but also the acceptance rate. In many applications, obtaining the joint posterior distribution of the quantity of interest over a subregion is sufficient; for example, one could obtain the joint distribution of a water-related quantity over a catchment area and then compute the distribution of functions of interests (e.g., the average) as necessary.
To demonstrate this idea, we first considered an example in sufficiently low dimensions so that the ‘global’ simulation of all censored locations had a reasonable acceptance rate above $10^{-4}$. Specifically, we considered Scenario 1 in Table 1, where the censoring threshold is zero. Figure 7 compares N=1, [^0] N=1{,}000 samples of the north-west quarter (i.e., $0,0.5]\times[0.5,1.0]$) generated by the global and regional sampling schemes using VMET.
@@
For the regional sampling scheme, we sampled the TMVN distribution defined over the region $0,0.6]\times[0.4,1.0]$, which enveloped the region of our interest, conditional on all observed locations. For the global sampling scheme, we needed to sample from a $380$-dimensional TMVN distribution, whereas for the regional scheme, we only needed to sample from a $125$-dimensional TMVN distribution, resulting in a $50$ to $100$ times difference in acceptance rates. The two groups of samples were very similar except for minor differences in the tails, which can be attributed to the reduced number of censored locations used in the regional sampling scheme. Conditioning on more censored locations would decrease the conditional mean under our Matérn covariance kernel. We also compared the RMSE for predicting the censored data in the north-west quarter, where the global and regional sampling produced errors of $0.24$ and $0.25$, respectively. In contrast, the GP with the LOD-augmented data produced a prediction error of $0.62$ and the stochastic approximation of expectation maximization (SAEM), a state-of-the-art method for sampling TMVN distributions from Ordoñez et al. (2018), produced an error of $0.47$. For SAEM, we set the MC sample size to $500$ as in Ordoñez et al. (2018), resulting in a computation time of approximately two hours, which was significantly higher than for our proposed methods. The regional scheme had much higher accuracy than the competing methods and the minor accuracy loss compared with the global scheme, we believe, is a good trade-off for the gain in computational efficiency.
@@
Next, we considered an example in higher dimensions to demonstrate that the VMET method can be used in combination with the regional sampling scheme to improve posterior inference when there are censored GP responses. We again considered Scenario 1 from Table 1, but now with n=6,400n=6{,}400. The region of interest and the region over which we drew TMVN samples were the same as in the low-dimensional example above. We had a total of 3,2783{,}278 censored locations, with 1,0751{,}075 inside our sampling region. N=1, [^0] N=1{,}000 samples were drawn and used in combination with all observed responses to predict the responses at $500$ randomly selected testing locations in our region of interest (i.e., $0,0.5]\times[0.5,1.0]$). For comparison, we predicted with the GP using the LOD-augmented data; we also attempted a comparison to SAEM, but the approach was computationally infeasible for sampling the 1,0751{,}075-dimensional TMVN distribution due to extremely low acceptance rates. The LOD-augmented GP produced an RMSE of $0.45$, whereas the RMSE of our method was only $0.19$. While the global approach cannot be computed as a gold-standard comparison method in this high-dimensional example, note that the nugget variance is $0.03$, indicating a minimum prediction RMSE of $0.17$. This highlights the benefits of properly taking the censored locations into consideration for posterior inference. In this example, the acceptance rate was approximately $2\times 10^{-5}$ and the sampling took less than 44 minutes. Combined with the results from Section 7.1, the VMET algorithm outlined in Algorithm 1 enables scalable (linear-complexity) computation for the partially censored MVN model in terms of parameter estimation and posterior inference, significantly outperforming the LOD-augmented GP and the SAEM approaches.
@@ 7.3 Analysis of groundwater contamination
We considered groundwater tetrachloroethylene concentrations from the United States Geological Survey (USGS) to demonstrate that our proposed VMET method can be used in combination with the censored MVN model to enhance the modeling of partially censored data. Tetrachloroethylene can be used in dry cleaning, manufacturing other chemicals, and cleaning metals, but exposure to tetrachloroethylene is harmful and may cause skin and respiratory problems. The data were collected at irregular locations across the United States over the time period 2000 to 2022. In this section, the locations $\{\mathbf{s}_{i}\in\mathbb{R}^{3}\}$ are spatial-temporal coordinates consisting of longitudes, latitudes, and dates. Overall, there are n=24,701n=24{,}701 spatio-temporal responses, among which there are 20,73020{,}730 censored responses due to detection limits. The detection thresholds are known and vary across spatio-temporal locations. Environmental contaminants are usually modeled using MVN distributions after a logarithm transformation (Helsel, 1990). During data preprocessing, we first applied a log-transform to both observations and detection thresholds. Next, we normalized the transformed observations and the detection thresholds with the mean and standard deviation of the transformed observations. The spatial and temporal coordinates were linearly scaled into the unit hypercube in $\mathbb{R}^{3}$, with the two spatial coordinates scaled by the same constant. The dataset, including observations and detection thresholds, is randomly split into 80% training and 20% testing.
@@
We used the state of Texas as our region of interest for posterior inference. Specifically, our goal was to make predictions over a dense regular $0.13^{\circ}\times 0.11^{\circ}$ longitude-latitude grid covering Texas on the last day of the dataset’s temporal range. Based on the predictions, we produced a ‘heatmap’ for the groundwater tetrachloroethylene concentrations in Texas. Conditional on the training observations, we drew N=1, [^0] N=1{,}000 samples of the TMVN distribution defined over the $546$ censored spatio-temporal points in the training dataset that are located in Texas. Inference over the grid as well as the testing spatio-temporal points in Texas was made by kriging using observed responses and the TMVN samples drawn previously. For comparison, we also considered predictions made by a GP using the augmented data where censored responses were replaced by their LODs. The SAEM method was computationally infeasible for sampling from the $546$-dimensional TMVN distribution, and it was hence excluded from our comparison. Figure 8 compares the heatmaps produced by the partially censored MVN model and the GP model.
@@ 8 Conclusions
Acknowledgments
The authors were partially supported by National Science Foundation (NSF) Grants DMS–1654083 and DMS–1953005. Support of JC’s research was also provided by the University of Houston. Support for MK’s research was also provided by the Office of the Vice Chancellor for Research and Graduate Education at the University of Wisconsin–Madison with funding from the Wisconsin Alumni Research Foundation. We would like to thank Kyle Messier for providing the tetrachloroethylene data and for several helpful comments and discussions.
Disclosure Statement
The authors have no competing interests to declare.